# On the Identifiability of Tensor Ranks via Prior Predictive Matching

### v1

@@ I Introduction
Tensor factorizations have become indispensable tools for analyzing multi-way data across diverse contexts and fields such as neuroscience (Cong et al., 2015; Mosayebi and Hossein-Zadeh, 2020), bioinformatics and genomics (Alter et al., 2003), chemometrics (Bro and Kiers, 2003), computational social science (Schein et al., 2015), network science (Contisciani et al., 2022), recommender systems (Bhargava et al., 2015; Gopalan et al., 2015), supervised learning (Stoudenmire and Schwab, 2016), and compression of large neural networks (Novikov et al., 2015). These methods decompose a high-dimensional tensor into a set of lower-dimensional factors, revealing latent structures and facilitating data interpretation and prediction (Chu and Ghahramani, 2009; Kolda and Bader, 2009; Cichocki et al., 2009).
@@
A fundamental challenge in the design and aplication of tensor models is the selection of their latent dimensions, i.e., the ranks ($r_{p}$), which dictate the model’s complexity and its ability to capture the observed statistics of the data, without overfitting. In contrast to simple matrix factorizations, most tensor models have multiple ranks, i.e., $p\geq 1$, depending on the specific form of the factorization. The determination of the tensor ranks is, in general, a NP-Complete problem for tensors over $\mathbb{Q}$ (Håstad, 1990), NP-Hard for tensors over $\mathbb{R}$ (Hillar and Lim, 2013) and hard to approximate (Swernofsky, 2018).
@@
In practice, Bayesian approaches to tensor factorization (Chu and Ghahramani, 2009; Zhao et al., 2015; Rai et al., 2014) provide a principled alternative to exhaustive search, enabling uncertainty quantification, prior incorporation, and even automatic rank inference via non-parametric methods (Porteous et al., 2008) or automatic relevance determination (Zhao et al., 2015). Yet, parametric models often demand user-specified ranks or strong priors, complicating deployment in real-world scenarios where domain knowledge is scarce.
@@
For matrix factorization models, the prior specification hurdle can be mitigated through prior predictive checks or empirical Bayes techniques (da Silva et al., 2023; Wang and Stephens, 2021). That is, we can determine the ideal rank by comparing observed data with data simulated from the model under alternative ranks. This established technique for matrix factorization inspires an analogous development for tensor factorization.
@@
Given this context, the fundamental question we ask here is when can a prior predictive moment matching approach identify the ranks of tensor factorization models. To answer the question, we need to establish a range of tools for connecting the moments of the prior predictive distribution with the ranks of the tensor model and an approach to use them for characterizing the rank identifiability of the model.
@@
In this work, we formulate a strategy for analyzing the identifiability of ranks in probabilistic tensor factorization models, by analyzing a log-linear system of equations derived from matching model-specific prior predictive moments with low-order moments of the observed tensor-data. We show that the algebraic structure of this system, which is determined by the chosen tensor decomposition, dictates whether the ranks can be uniquely identified by the method of moment-matching. We formalize this analysis by transforming the multiplicative moment equations into a log-linear system, where identifiability is equivalent to the solvability of the system with respect to the unknown ranks. Notably, we provide a clear theoretical characterization and distinction between different tensor models, demonstrating that rank identifiability is a direct consequence of the interaction topology of their latent factors. Besides, we showcase how ranks can be empirically estimated from observed tensor-data according to the prior predictive matching framework.
@@
    We apply this framework to four foundational tensor decompositions: Tucker (Tucker, 1966),
    PARAFAC/CP (Harshman et al., 1970; Carroll and Chang, 1970), Tensor Train ($T$, Oseledets, 2011), and Tensor Ring (TR, Zhao et al., 2016).
@@
Table 1: Summary of identifiability and closed-form estimators from second moments. All $v_{S}$ are pure interaction terms obtained via exact-sharing covariances and Möbius inversion; $E[Y]=\mathbb{E}[Y]$.
Model Identifiable? Rank parameters Closed-form from observables
Tucker No (from 2nd moments) $(r_{1},\dots,r_{M})$ Symmetry induces $M$ identities $\log v_{\{p\}}-2\log\mathbb{E}[Y]=\log v_{\{G,p\}}-\log v_{\{G\}}$; the reduced log-rank system is degenerate (non-identifiable).
PARAFAC/CP Yes $r$ For any $p\neq q$:  $r=\dfrac{v_{\{p,q\}}\,(\mathbb{E}[Y])^{2}}{v_{\{p\}}\,v_{\{q\}}}$.
Tensor Train (TT) Yes $(r_{1},\dots,r_{M-1})$ {rp=v{p,p+1}v{p-1,p+2}v{p+1}v{p-1,p,p+2} Interior 1<p<M-1r1=v{1,2}v{3}v{2}v{1,3}  p=1 and M ≥3rM-1=v{M-1,M}v{M-3}v{M}v{M-3,M-1}  p=M-1 and M>3
Tensor Ring (TR) Yes $(r_{1},\dots,r_{M})$ Define $\xi_{p}=\dfrac{(\mathbb{E}[Y])^{2}v_{\{p,p+1\}}}{v_{\{p\}}v_{\{p+1\}}}=r_{p}\tfrac{r_{p-1}}{r_{p+1}}$. Solve $x_{p}+x_{p-1}-x_{p+1}=\psi_{p}$ with $x_{p}=\log r_{p}$, $\psi_{p}=\log\xi_{p}$; circulant system is invertible for $M\geq 3$.
II A General Framework for Rank Identifiability
@@
Let $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ be an observed tensor with $M$ modes. A mode of an order-$M$ tensor is one of its $M$ dimensions. Let $\mathbf{i_{\alpha}}=(i_{m})_{m\in[M]}\in\bigtimes_{m=1}^{M}[N_{m}]$ be the multi-index from the observed tensor domain, where $N_{m}]=\{1,\dots,N_{m}\}$ . Let $\mathbf{r}=(r_{p})_{p\in[R]}$ be the multi-rank vector of size $R\leq M$. Let $\mathbf{i_{\beta}}=(i_{1},\dots,i_{R})\in\bigtimes_{p=1}^{R}[r_{p}]$ be the multi-index of the product of rank domains $r_{p}]$. For multi-indices $\mathbf{i},\mathbf{j}$, the shared-mode set is $S(\mathbf{i},\mathbf{j})=\{k:i_{k}=j_{k}\}$. We use bold sans-serif for tensors ($\bm{\eta}$) and vectors/matrices ($\bm{\theta}^{(p)}$). All latent factor multi-indices $\mathbf{i}_{p}(\alpha,\beta)$ are formed by combining some components of $\mathbf{i_{\alpha}}$ and some components of $\mathbf{i_{\beta}}$, and each factor is probabilistically defined as $\theta^{(p)}_{\mathbf{i}_{p}(\alpha,\beta)}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, i.i.d. across indices with $p$ indexing and grouping the factors.
Tensor diagram.
We use compact tensor networks to visualize factorizations. Each node is a tensor (core/factor); solid legs are indices. A single dangling leg denotes a physical mode (observed dimension); edges connecting nodes are latent “bond” indices whose sizes are the ranks. Chains of nodes depict TT; a cycle depicts TR; a star-like spoke pattern with a central core depicts Tucker; CP is a sum of rank-1 outer products (often drawn as $r$ parallel spoke motifs). Diagrammatic equalities correspond to index contractions; our counting arguments track how many independent latent sums (i.e., free edges) remain under each exact-sharing pattern, which become exponents of the ranks in the monomials.
Definition II.1 (The Probabilistic Tensor-Model).
@@
We assume that elements in $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ are probabilistically generated from a rate tensor $\bm{\eta}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$, whose structure is defined by a (model-dependent) tensor decomposition with latent factors of ranks $\mathbf{r}=(r_{1},\ldots,r_{R})$ of the following generic form:
$\eta_{\mathbf{i_{\alpha}}}=\sum_{\mathbf{i_{\beta}}}\prod_{p}\theta^{(p)}_{\mathbf{i}_{p}(\alpha,\beta)},$ (1)
@@
with latent indices defined by $\mathbf{i}_{p}(\alpha,\beta)$ and parameters $\theta^{(p)}_{\mathbf{i}_{\{\alpha,\beta\}}}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, drawn independently from a factor $p$-specific location-scale prior distribution $\pi_{p}(\cdot)$ with up-to-second order moments $(\mu_{p},\sigma_{p}^{2})$.
@@
Tensor observations $Y_{i}$ are conditionally independent given the rate tensor $\bm{\eta}$, with conditional moments $\mathbb{E}[Y_{i}|\bm{\eta}]=\eta_{i}$ and $\mathrm{Var}(Y_{i}|\bm{\eta})=\phi(\eta_{i})$ for some model-dependent positive function $\phi(\cdot)$.
@@
This formulation allows for different observation models (e.g. Poisson with $\phi(\eta)=\eta$, or Gaussian with $\phi(\eta)=\sigma_{Y}^{2}$ variance), and it generalizes existing matrix factorization frameworks (da Silva et al., 2023), setting the stage for the prior predictive moment matching analysis. Commonly used factorization models such as Tucker, PARAFAC/CP, TR and TT, are all instances of this general formulation  [^1] If $\mathbf{i}_{p}(\alpha,\beta)$ is formed by taking the $p$-th components of $\mathbf{i}_{\alpha}$ and $\mathbf{i}_{\beta}$ respectively, for $p>0$, and for $p=0$ define a special $\mathbf{i}_{0}(\alpha,\beta)=\mathbf{i}_{\beta}$, we obtain Tucker. While doing the same but defining $\mathbf{i}_{\beta}$ as a one dimensional index shared by all, without a special $p=0$ index, we obtain PARAFAC/CP..
@@
The second moments of the observable tensor-data can be described by two key quantities: the total covariance and pure interaction terms, which form the key building blocks for the theoretical analysis that follows.
Definition II.2 (Total Covariance and Pure Interaction Terms).
Let $\mathbf{i}=(i_{1},\dots,i_{M})$ and $\mathbf{j}=(j_{1},\dots,j_{M})$ be two multi-indices from the tensor’s domain. Let $S(\mathbf{i},\mathbf{j})=\{k\mid i_{k}=j_{k}\}$ be the set of indices they share. The total observable covariance, $C_{S}$, is the theoretical covariance between two multi-index tensor entries $\mathbf{i}$ and $\mathbf{j}$:
@@
Connecting theorical and observable moments.
@@
Proposition II.3 (Polynomial Structure of Rate Moments).
For a probabilistic tensor model congruent with Definition II.1, every pure interaction term $v_{S}$ is a monomial in the unknown model hyperparameters ($r_{p},\mu_{p},\sigma^{2}_{p}$). Therefore, the marginal mean $\mathbb{E}[\bm{\eta}]$, the variance $\mathrm{Var}(\bm{\eta})$, and the covariance terms $\mathrm{Cov}(\eta_{\mathbf{i}},\eta_{\mathbf{j}})$ can be expressed as polynomials of $\{r_{p},\mu_{p},\sigma_{p}^{2}\}$.
@@
Each rate element $\eta_{\mathbf{i}_{\alpha}}$ is a sum-product of the latent factor elements $\{\theta^{(p)}_{\mathbf{i}_{\{\alpha,\beta\}}}\}$. Applying total expectation we obtain $\mathbb{E}[\bm{Y}_{\mathbf{i_{a}}}]=\sum_{i_{\beta}}\underset{p:i_{p}(\alpha,\beta)}{\prod}\mu_{p}$, with the summation index depending on the rank dimensions, and resulting in the monomial given by products of ranks $\{r_{p}\}$ and means $\{\mu_{p}\}$. Applying total covariance and conditional independence, the computation leads to moments $\mathbb{E}[\eta_{\mathbf{i_{\alpha}}}\eta_{\mathbf{j}_{\alpha}}]$, which resolve into a double sum-product over latent indexes, with $\sigma_{p}^{2}$ appearing for factors with the same latent index, $\mu_{p}^{2}$ for independent factors, and powers of $r_{p}$ from interactions of repeated sums and multiplications, leading to a polynomial on prior parameters $\{\mu_{p},\sigma_{p}^{2}\}$ and ranks $\{r_{p}\}$. ∎
II.1 The Log-Linear System of Prior Predictive Moment Equations
The insight established by Proposition II.3 is that observable moments (like covariances $C_{S}$) are polynomials of the model parameters. More precisely, the individual terms of these polynomials are the pure interaction terms $v_{S}$, which are monomials. This monomial (multiplicative) structure is ideal for linearization. Namely, by taking the logarithm of the equation for each pure interaction term, we transform the system of prior predictive moments into a linear one, amenable to standard algebraic analysis.
Definition II.4 (The Log-Linear System).
@@ II.1 The Log-Linear System of Prior Predictive Moment Equations
Proposition II.5 (Principle of Rank Identifiability).
The ranks $\{r_{p}\}$ of a tensor model are identifiable from the second moments of tensor-data if and only if the unknown prior-moment parameters can be algebraically eliminated from the full log-linear system of moments $\mathbf{A}\mathbf{x}=\mathbf{b}$, to yield a non-trivial reduced system of the log-ranks $\mathbf{x}_{r}$ alone: i.e., $\mathbf{A}_{\text{red}}\mathbf{x}_{r}=\mathbf{b}_{\text{red}}$, where the reduced design matrix $\mathbf{A}_{\text{red}}$ has full rank. A full-rank $\mathbf{A}_{\text{red}}$ guarantees a unique solution for the ranks, meaning they occupy a parameter subspace that can be uniquely determined from the data’s moments.
Propositions II.3 and II.5 establish a direct link between the model-dependent, prior predictive moments and the moments of the observed tensor-data. This connection is formalized through an identifiability analysis of a reduced, rank-specific log-linear system. However, the utility of this link for rank estimation is fully model-dependent.
We first present in Section III a cautionary example where the dependencies between prior hyperparameters and observed second moments are not identifiable for a popular tensor model. In contrast, we highlight in Section IV several models where identifiability holds, enabling the design and implementation of a novel method for estimating tensor ranks directly and exclusively from observed moments.
III A Cautionary Tale: The Tucker Model

### v2

@@ 1 INTRODUCTION
Tensor factorizations have become indispensable tools for analyzing multi-way data across diverse contexts and fields such as neuroscience (Cong et al., 2015; Mosayebi and Hossein-Zadeh, 2020), bioinformatics and genomics (Alter et al., 2003), chemometrics (Bro and Kiers, 2003), computational social science (Schein et al., 2015), network science (Contisciani et al., 2022), recommender systems (Bhargava et al., 2015; Gopalan et al., 2015), supervised learning (Stoudenmire and Schwab, 2016), compression of large neural networks (Novikov et al., 2015), and trace back to early work by Hitchcock (1927). These methods decompose a high-dimensional tensor into a set of lower-dimensional factors, revealing latent structures and facilitating data interpretation and prediction (Chu and Ghahramani, 2009; Kolda and Bader, 2009; Cichocki et al., 2009).
@@
A fundamental challenge in the design and application of tensor models is the selection of their latent dimensions, i.e., the ranks ($r_{p}$), which dictate the model’s complexity and its ability to capture the observed statistics of the data, without overfitting. In contrast to simple matrix factorizations, most tensor models have multiple ranks, i.e., $p\geq 1$, depending on the specific form of the factorization. The determination of the tensor ranks is, in general, a NP-Complete problem for tensors over $\mathbb{Q}$ (Håstad, 1990), NP-Hard for tensors over $\mathbb{R}$ (Hillar and Lim, 2013) and hard to approximate in general (Swernofsky, 2018).
@@
Bayesian approaches to tensor factorization (Chu and Ghahramani, 2009; Zhao et al., 2015; Rai et al., 2014) provide a principled alternative to exhaustive search, enabling uncertainty quantification, prior incorporation, and even automatic rank inference via non-parametric methods (Porteous et al., 2008) or automatic relevance determination (Zhao et al., 2015).
@@
However, parametric models often demand user-specified ranks or strong priors, complicating deployment in real-world scenarios where domain knowledge is scarce. For matrix factorization models, the prior specification hurdle can be mitigated through prior predictive checks or empirical Bayes techniques (da Silva et al., 2023; Wang and Stephens, 2021). That is, we can determine the matrix rank by comparing observed data with data simulated from the model under alternative ranks. This established technique for matrix factorization inspires an analogous development for tensor factorization we study here.
@@
The fundamental question we ask here is when can a prior predictive moment matching approach identify the ranks of tensor factorization models. To answer the question, we need to establish a range of tools for connecting the moments of the prior predictive distribution with the ranks of the tensor model and an approach to use them for characterizing the rank identifiability of the model.
@@
In this work, we formulate a strategy for analyzing the identifiability of ranks in probabilistic tensor factorization models, by analyzing a log-linear system of equations derived from matching model-specific prior predictive moments with low-order moments of the observed tensor-data. We show that the algebraic structure of this system, which is determined by the chosen tensor decomposition, dictates whether the ranks can be uniquely identified by the method of moment-matching. We formalize this analysis by transforming the moment equations into a log-linear system, where identifiability is equivalent to the solvability of the system with respect to the unknown ranks.
Fundamentally, this framework establishes a link between a tensor factorization model’s structure and its rank identifiability. We provide a clear theoretical characterization and distinction between different tensor models, demonstrating that rank identifiability is a direct consequence of the interaction topology of their latent factors. In addition, we showcase how ranks can be empirically estimated from observed tensor-data according to the prior predictive matching method.
@@
    We apply this framework to four foundational tensor decompositions: Tucker (Tucker, 1966), PARAFAC/CP (Hitchcock, 1927; Harshman and others, 1970; Carroll and Chang, 1970), Tensor Train ($T$, Oseledets, 2011), and Tensor Ring (TR, Zhao et al., 2016).
@@
2 A GENERAL FRAMEWORK FOR RANK IDENTIFIABILITY
@@
Let $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ be an observed tensor with $M$ modes. A mode of an order-$M$ tensor is one of its $M$ dimensions. Let $\mathbf{i}=(i_{m})_{m\in[M]}\in\bigtimes_{m=1}^{M}[N_{m}]$ be the multi-index from the observed tensor domain, where $N_{m}]=\{1,\dots,N_{m}\}$ . Let $\mathbf{r}=(r_{p})_{p\in[R]}$ be the multi-rank vector of size $R\leq M$. Let $\beta\in\mathcal{K}=\bigtimes_{p=1}^{R}[r_{p}]$ be the latent multi-index over the rank domains. For multi-indices $\mathbf{i},\mathbf{j}$, the shared-mode set is $S(\mathbf{i},\mathbf{j})=\{k:i_{k}=j_{k}\}$. We use bold sans-serif for tensors ($\bm{\eta}$) and vectors/matrices ($\bm{\theta}^{(p)}$). All latent factor multi-indices $\mathbf{i}_{p}(\mathbf{i},\beta)$ are formed by combining some components of $\mathbf{i}$ and the latent index $\beta$, and each factor is probabilistically defined as $\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, i.i.d. across indices with $p\in\mathcal{P}$ indexing and grouping the factors.
Definition 2.1 (The Probabilistic Tensor-Model).
@@
We assume that elements in tensor $\bm{Y}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$ are probabilistically generated from a rate tensor $\bm{\eta}\in\mathbb{R}^{N_{1}\times\dots\times N_{M}}$, whose structure is defined by a (model-dependent) tensor decomposition with latent factors of ranks $\mathbf{r}=(r_{1},\ldots,r_{R})$ of the following generic form:
$\eta_{\mathbf{i}}=\sum_{\beta\in\mathcal{K}}\prod_{p\in\mathcal{P}}\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)},$ (1)
@@
with latent indices defined by $\mathbf{i}_{p}(\mathbf{i},\beta)$ and parameters $\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)}\sim\pi_{p}(\mu_{p},\sigma_{p}^{2})$, drawn independently from a factor $p$-specific location-scale prior distribution $\pi_{p}(\cdot)$ with up-to-second order moments $(\mu_{p},\sigma_{p}^{2})$.
@@
Tensor observations $Y_{\mathbf{i}}$ are conditionally independent given the rate tensor $\bm{\eta}$, with conditional moments $\mathbb{E}[Y_{\mathbf{i}}|\bm{\eta}]=\eta_{\mathbf{i}}$ and $\mathrm{Var}(Y_{\mathbf{i}}|\bm{\eta})=\phi(\eta_{\mathbf{i}})$ for some model-dependent positive function $\phi(\cdot)$.
@@
This probabilistic tensor formulation allows for different observation models (e.g. Poisson with $\phi(\eta)=\eta$, or Gaussian with $\phi(\eta)=\sigma_{Y}^{2}$ variance), and it generalizes existing matrix factorization frameworks (da Silva et al., 2023), setting the stage for the prior predictive moment matching analysis. Commonly used factorization models such as Tucker, PARAFAC/CP, TR and TT, are all instances of this general formulation  [^1] If $\mathbf{i}_{p}(\mathbf{i},\beta)$ is formed by taking the $p$-th components of $\mathbf{i}$ and $\beta$ respectively, for $p>0$, and for $p=0$ define a special $\mathbf{i}_{0}(\mathbf{i},\beta)=\beta$, we obtain Tucker. While doing the same but defining $\beta$ as a one dimensional index shared by all, without a special $p=0$ index, we obtain PARAFAC/CP..
@@
The second moments of the observable tensor-data can be described by two key quantities: the total covariance and the pure interaction terms, which form the key building blocks for the analysis that follows.
Definition 2.2 (Total Covariance and Pure Interaction Terms).
Let $\mathbf{i}=(i_{1},\dots,i_{M})$ and $\mathbf{j}=(j_{1},\dots,j_{M})$ be two multi-indices from the tensor’s domain. Let $S(\mathbf{i},\mathbf{j})=\{k\mid i_{k}=j_{k}\}$ be the set of indices they share.
The total observable covariance, $C_{S}$, is the theoretical covariance between two multi-index tensor entries $\mathbf{i}$ and $\mathbf{j}$:
@@
Connecting theoretical and observable moments.
@@
Proposition 2.3 (Polynomial Structure of Rate Moments).
For a probabilistic tensor model congruent with Definition 2.1, every pure interaction term $v_{S}$ is a monomial in the unknown model hyperparameters ($r_{p},\mu_{p},\sigma^{2}_{p}$). Therefore, the marginal mean $\mathbb{E}[\bm{\eta}]$, the variance $\mathrm{Var}(\bm{\eta})$, and the covariance terms $\mathrm{Cov}(\eta_{\mathbf{i}},\eta_{\mathbf{j}})$ can be expressed as polynomials of $\{r_{p},\mu_{p},\sigma_{p}^{2}\}$.
@@
Each rate element $\eta_{\mathbf{i}}$ is a sum-product of the latent factor elements $\{\theta^{(p)}_{\mathbf{i}_{p}(\mathbf{i},\beta)}\}$. Applying total expectation we obtain $\mathbb{E}[\bm{Y}_{\mathbf{i}}]=\sum_{\beta\in\mathcal{K}}\underset{p}{\prod}\mu_{p}$, with the summation index depending on the rank dimensions, and resulting in the monomial given by products of ranks $\{r_{p}\}$ and means $\{\mu_{p}\}$. Applying total covariance and conditional independence, the computation leads to moments $\mathbb{E}[\eta_{\mathbf{i}}\eta_{\mathbf{j}}]$, which resolve into a double sum-product over latent indexes, with $\sigma_{p}^{2}$ appearing for factors with the same latent index, $\mu_{p}^{2}$ for independent factors, and powers of $r_{p}$ from interactions of repeated sums and multiplications, leading to a polynomial on prior parameters $\{\mu_{p},\sigma_{p}^{2}\}$ and ranks $\{r_{p}\}$. ∎
2.1 The Log-Linear System of Prior Predictive Moment Equations
The insight unveiled by Proposition 2.3 is that observable moments (like covariances $C_{S}$) are polynomials of the model parameters. More precisely, the individual terms of these polynomials are the pure interaction terms $v_{S}$, which are monomials. This monomial (multiplicative) structure is ideal for linearization. Namely, by taking the logarithm of the equation for each pure interaction term, we transform the system of prior predictive moments into a linear one, amenable to standard algebraic analysis.
Definition 2.4 (The Log-Linear System of Rate Moments).
@@ 2.1 The Log-Linear System of Prior Predictive Moment Equations
Proposition 2.5 (Principle of Rank Identifiability).
The ranks $\{r_{p}\}$ of a tensor model are identifiable from the second moments of tensor-data if and only if the unknown prior-moment parameters can be algebraically eliminated from the full log-linear system of moments $\mathbf{A}\mathbf{x}=\mathbf{b}$.
Consequently, it yields a non-trivial reduced system of the log-ranks $\mathbf{x}_{r}$ alone: i.e., $\mathbf{A}_{\text{red}}\mathbf{x}_{r}=\mathbf{b}_{\text{red}}$, where the reduced design matrix $\mathbf{A}_{\text{red}}$ has full rank. A full-rank $\mathbf{A}_{\text{red}}$ guarantees a unique solution for the ranks, meaning they occupy a parameter subspace that can be uniquely determined from the data’s moments.
Propositions 2.3 and 2.5 establish a direct link between the model-dependent, prior predictive moments and the moments of the observed tensor-data. This connection is formalized through an identifiability analysis of a reduced, rank-specific log-linear system.
However, we note that the utility of this link for rank estimation is fully model-dependent. We first present in Section 3 a cautionary example where the dependencies between prior hyperparameters and observed second moments are not identifiable for the popular Tucker tensor model. In contrast, we highlight in Section 4 several models where identifiability holds, enabling the design and implementation of a novel method for estimating tensor ranks directly and exclusively from observed moments.
3 A CAUTIONARY TALE: THE TUCKER MODEL